# LLM4Decompile: Decompiling Binary Code with Large Language Models

### v2

@@ Abstract
Decompilation aims to convert binary code to high-level source code, but traditional tools like Ghidra often produce results that are difficult to read and execute. Motivated by the advancements in Large Language Models (LLMs), we propose LLM4Decompile, the first and largest open-source LLM series (1.3B to 33B) trained to decompile binary code. We optimize the LLM training process and introduce the LLM4Decompile-End models to decompile binary directly. The resulting models significantly outperform GPT-4o and Ghidra on the HumanEval and ExeBench benchmarks by over 100%. Additionally, we improve the standard refinement approach to fine-tune the LLM4Decompile-Ref models, enabling them to effectively refine the decompiled code from Ghidra and achieve a further 16.2% improvement over the LLM4Decompile-End. LLM4Decompile [^1] https://github.com/albertan017/LLM4Decompile demonstrates the potential of LLMs to revolutionize binary code decompilation, delivering remarkable improvements in readability and executability while complementing conventional tools for optimal results.
@@
Hanzhuo Tan, Qi Luo, Jing Li, Yuqun Zhang Southern University of Science and Technology The Hong Kong Polytechnic University
@@ 1 Introduction
Decompilation, the reverse process of converting machine code or binary code into a high-level programming language, facilitates various reverse engineering tasks such as vulnerability identification, malware research, and legacy software migration Brumley et al. (2013); Katz et al. (2018); Hosseini and Dolan-Gavitt (2022); Xu et al. (2023); Armengol-Estapé et al. (2023); Jiang et al. (2023); Wong et al. (2023); Hu et al. (2024). Decompilation is challenging due to the loss of information inherent in the compilation process, particularly finer details such as variable names Lacomis et al. (2019) and fundamental structures like loops and conditionals Wei et al. (2007). To address these challenges, numerous tools have been developed for decompilation, with Ghidra Ghidra (2024) and IDA Pro Hex-Rays (2024) being the most commonly used. Although these tools have the capability to revert binary code to high-level pseudo-code, the outputs often lack readability and re-executability Liu and Wang (2020a); Wang et al. (2017), which are essential for applications like legacy software migration and security instrumentation tasks Wong et al. (2023); Dinesh et al. (2020).
@@
Figure 1 illustrates the transformation from the source C code to a binary file, assembly code (ASM), and pseudo-code decompiled from Ghidra. In this pseudo-code, the original nested for structure is replaced with a less intuitive combination of a do-while loop inside another while loop. Furthermore, array indexing like num[i] is decompiled into complicated pointer arithmetic such as *(float *)(param_2 + (long)local_24 * 4). The decompiled output also exhibits syntactical errors, with the function return type being converted to undefined4. Overall, traditional decompilation tools often strip away the syntactic clarity provided by high-level languages and do not ensure the correctness of syntax, posing significant challenges even for skilled developers to reconstruct the algorithmic logic Wong et al. (2023); Hu et al. (2024)
@@
To address the limitations of previous studies, we propose LLM4Decompile, the first and largest open-source LLM series with sizes ranging from 1.3B to 33B parameters specifically trained to decompile binary code. To the best of our knowledge, there’s no previous study attempts to improve the capability of LLM-based decompilation in such depth or incorporate such large-scale LLMs. Based on the End2end-Decompile approach, we introduce three critical steps: data augmentation, data cleaning, and two-stage training, to optimize the LLM training process and introduce the LLM4Decompile-End models to decompile binary directly. Specifically, our LLM4Decompile-End-6.7B model demonstrates a successful decompilation rate of 45.4% on HumanEval Chen et al. (2021) and 18.0% on ExeBench Armengol-Estapé et al. (2022), far exceeding Ghidra or GPT-4o by over 100%. Additionally, we improve the Refined-Decompile strategy by examining the efficiency of Ghidra’s decompilation process, augmenting and filtering data to fine-tune the LLM4Decompile-Ref models, which excel at refining Ghidra’s output. Experiments suggest a higher performance ceiling for the enhanced Refined-Decompile approach, with 16.2% improvement over LLM4Decompile-End. Additionally, we assess the risks associated with the potential misuse of our model under obfuscation conditions commonly used in software protection. Our findings indicate that neither our approach nor Ghidra can effectively decompile obfuscated code, mitigating concerns about unauthorized use for infringement of intellectual property.
@@
  • We optimize the LLM training process and introduce LLM4Decompile-End models, which set a new performance standard of direct binary decompilation, significantly surpassing GPT-4o and Ghidra by over 100% on the HumanEval and ExeBench benchmarks.
  • We improve the Refined-Decompile approach to fine-tune the LLM4Decompile-Ref models, enabling them to effectively refine the decompiled results from Ghidra and achieve further 16.2% enhancements over LLM4Decompile-End.
@@ 2 Related Work
The practice of reversing executable binaries to their source code form, known as decompilation, has been researched for decades Miecznikowski and Hendren (2002); Nolan (2012); Katz et al. (2019). Traditional decompilation relies on analyzing the control and data flows of program Brumley et al. (2013), and employing pattern matching, as seen in tools like Hex-Rays Ida pro Hex-Rays (2024) and Ghidra Ghidra (2024). These systems attempt to identify patterns within a program’s control-flow graph (CFG) that corresponding to standard programming constructs such as conditional statements or loops. However, the output from such decompilation processes tends to be a source-code-like representation of assembly code, including direct translations of variables to registers, use of gotos, and other low-level operations instead of the original high-level language constructs. This output, while often functionally similar to the original code, is difficult to understand and may not be re-executable Liu and Wang (2020b); Wong et al. (2023). Drawing inspiration from neural machine translation, researchers have reformulated decompilation as a translation exercise, converting machine-level instructions into readable source code Katz et al. (2019). Initial attempts in this area utilized recurrent neural networks (RNNs) Katz et al. (2018) for decompilation, complemented by error-correction techniques to enhance the outcomes.
@@
Motivated by the success of Large Language Models Li et al. (2023); Rozière et al. (2023); Guo et al. (2024), researchers have employed LLMs for decompilation, primarily through two approaches—Refined-Decompile and End2end-Decompile. In particular, Refined-Decompile prompts the LLMs to refine results from traditional decompilation tools like Ghidra or IDA Pro. For instance, DeGPT Hu et al. (2024) enhances Ghidra’s readability by reducing cognitive load by 24.4%, while DecGPT Wong et al. (2023) increases IDA Pro’s re-executability rate to over 75% by integrating error messages into its refinement process. These approaches, however, largely ignore the fact that LLMs are designed primarily for high-level programming languages Li et al. (2023); Rozière et al. (2023); Guo et al. (2024), and their effectiveness with binary files is not well-established. End2end-Decompile, on the other hand, fine-tunes LLMs to decompile binaries directly. Early open-source models like BTC Hosseini and Dolan-Gavitt (2022) and recent development Slade Armengol-Estapé et al. (2023) adopt the language model with around 200 million parameters Lewis et al. (2020) to fine-tune for decompilation. While Nova Jiang et al. (2023), which is not open-sourced, develops a binary LLM with 1 billion parameters and fine-tunes it for decompilation. Consequently, the largest open-source model in this domain is limited to 200M. Whereas utilizing larger models trained on broader datasets has proven to substantially improve the performance Hoffmann et al. (2024); Kaplan et al. (2020); Rozière et al. (2023).
@@ 3.1.2 Optimize LLM4Decompile-End
Our final step for training optimization aims to educate the model with binary knowledge, and includes two-stage training. In the first stage, we train the model with a large corpus of compilable but not linkable (executable) data. Note that it’s significantly easier to extract C code that is compilable but not linkable da Silva et al. (2021); Armengol-Estapé et al. (2022). Such not-executable binary object code will closely resemble its executable version except it lacks linked addresses for external symbols. Therefore, in the first stage, we use the extensive compilable codes to ground our model in binary knowledge. In the second stage, we refine the model using executable code to ensure its practical applicability. We also conduct an ablation study for the two-stage training in Section 4.1.2.
@@ 3.2.2 Refine Ghidra by LLM4Decompile-Ref
Decompiling the executable code with Ghidra (Figure 3) is time-consuming due to the complex nature of the executables in ExeBench, which include numerous external functions and IO wrappers. Ghidra Headless requires 2 seconds per sample using 128-core multiprocessing. Given such a high computational load, and the high similarities between non-executable and executable binaries, we choose to decompile the non-executable files using Ghidra. This choice significantly reduces the time to 0.2 seconds per sample, enabling us to efficiently gather large amounts of training data.
@@ 4.1.1 Experimental Setups
To evaluate the models, we introduce HumanEval Chen et al. (2021) and ExeBench Armengol-Estapé et al. (2022) benchmarks. HumanEval is the leading benchmark for code generation assessment and includes 164 programming challenges with accompanying Python solutions and assertions. We converted these Python solutions and assertions into C, making sure that they can be compiled with the GCC compiler using standard C libraries and pass all the assertions, and name it HumanEval-Decompile. ExeBench consists of 5000 real-world C functions taken from GitHub with IO examples. Note that the HumanEval-Decompile consists of individual functions that depend only on the standard C library. In contrast, ExeBench includes functions extracted from real-world projects with user-defined structures and functions.
@@
Model/Benchmark HumanEval-Decompile ExeBench O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG DeepSeek-Coder-6.7B 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 0.0000 GPT-4o 0.3049 0.1159 0.1037 0.1159 0.1601 0.0443 0.0328 0.0397 0.0343 0.0378 LLM4Decompile-End-1.3B 0.4720 0.2061 0.2122 0.2024 0.2732 0.1786 0.1362 0.1320 0.1328 0.1449 LLM4Decompile-End-6.7B 0.6805 0.3951 0.3671 0.3720 0.4537 0.2289 0.1660 0.1618 0.1625 0.1798 LLM4Decompile-End-33B 0.5168 0.2556 0.2415 0.2475 0.3154 0.1886 0.1465 0.1396 0.1411 0.1540
@@
Model/Benchmark HumanEval-Decompile ExeBench O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG Compilable-1.3B 0.4268 0.1646 0.1646 0.1707 0.2317 0.0568 0.0446 0.0416 0.0443 0.0468 Compilable-6.7B 0.5183 0.3354 0.3232 0.3232 0.3750 0.0752 0.0649 0.0671 0.0660 0.0683 Executable-1.3B 0.1951 0.1280 0.1280 0.1159 0.1418 0.2194 0.1946 0.1931 0.1950 0.2005 Executable-6.7B 0.3720 0.1829 0.2256 0.1707 0.2378 0.2938 0.2598 0.2591 0.2549 0.2669
@@
We selected two key baselines for comparison. First, GPT-4o OpenAI (2023) represents the most capable LLMs, providing an upper bound on LLM performance. Second, DeepSeek-Coder Guo et al. (2024) is selected as the current SOTA open-source Code LLM. It represents the forefront of publicly available models specifically tailored for coding tasks. While recent work Slade Armengol-Estapé et al. (2023) fine-tunes language model for decompilation, it relies on intermediate compiler outputs, specifically, the *.s files. In practice, however, such intermediate files are rarely released by developers. Therefore, we focus on a more realistic approach, and consider decompilation only from the binaries, for further discussions please refer to Appendix A.
@@ 4.1.2 Experimental Results
Table 3: Main comparison of Refined-Decompile approaches for re-executability rate and Edit Similarity on HumanEval-Decompile benchmark. “+GPT-4o” refers to enhance the Ghidra results with GPT-4o, “+LLM4Decompile-Ref” means refining Ghidra results with the fine-tuned LLM4Decompile-Ref models.
@@
As discussed in Section 4.1.1, our training data comprises two distinct sets: 7.2 million compilable functions (non-executable) and 1.6M executable functions. We conducted an ablation study using these datasets, and the results are displayed in Table 2. Here, “Compilable” denotes the model trained solely on compilable data, while “Executable” indicates models trained exclusively on executable data. Notably, the binary object from compilable functions lacks links to function calls, which is similar in text distribution to the HumanEval-Decompile data, consisting of single functions dependent only on standard C libraries. Consequently, the 6.7B model trained only on compilable data successfully decompiled 37.5% of HumanEval-Decompile functions, but only 6.8% on ExeBench, which features real functions with extensive user-defined functions. On the other hand, the 6.7B model trained solely on executable data achieved a 26.7% re-executability rate on the ExeBench test set but faced challenges with single functions, with only a 23.8% success rate on HumanEval-Decompile due to the smaller size of the training corpus. Limited by the space, we present further analysis in Appendix B.

### v3

@@ Abstract
Decompilation aims to convert binary code to high-level source code, but traditional tools like Ghidra often produce results that are difficult to read and execute. Motivated by the advancements in Large Language Models (LLMs), we propose LLM4Decompile, the first and largest open-source LLM series (1.3B to 33B) trained to decompile binary code. We optimize the LLM training process and introduce the LLM4Decompile-End models to decompile binary directly. The resulting models significantly outperform GPT-4o and Ghidra on the HumanEval and ExeBench benchmarks over 100% in terms of re-executability rate. Additionally, we improve the standard refinement approach to fine-tune the LLM4Decompile-Ref models, enabling them to effectively refine the decompiled code from Ghidra and achieve a further 16.2% improvement over the LLM4Decompile-End. LLM4Decompile [^1] https://github.com/albertan017/LLM4Decompile demonstrates the potential of LLMs to revolutionize binary code decompilation, delivering remarkable improvements in readability and executability while complementing conventional tools for optimal results.
@@
Hanzhuo Tan1,2, Qi Luo1, Jing Li2,3, Yuqun Zhang1thanks: Yuqun Zhang is the corresponding author., 1Department of Computer Science and Engineering, Southern University of Science and Technology, Shenzhen, China 2Department of Computing, The Hong Kong Polytechnic University, HKSAR, China 3 Research Centre for Data Science & Artificial Intelligence hanzhuo.tan@connect.polyu.hk, 12232440@mail.sustech.edu.cn, jing-amelia.li@polyu.edu.hk, zhangyq@sustech.edu.cn
@@ 1 Introduction
Decompilation, the reverse process of converting machine code or binary code into a high-level programming language, facilitates various reverse engineering tasks such as vulnerability identification, malware research, and legacy software migration Brumley et al. (2013); Katz et al. (2018); Hosseini and Dolan-Gavitt (2022); Xu et al. (2023); Armengol-Estapé et al. (2023); Jiang et al. (2023); Huaqingweiyang (2024); Wong et al. (2023); Hu et al. (2024). Decompilation is challenging due to the loss of information inherent in the compilation process, particularly finer details such as variable names Lacomis et al. (2019) and fundamental structures like loops and conditionals Wei et al. (2007). To address these challenges, numerous tools have been developed for decompilation, with Ghidra Ghidra (2024a) and IDA Pro Hex-Rays (2024) being the most commonly used. Although these tools have the capability to revert binary code to high-level pseudo-code, the outputs often lack readability and re-executability Liu and Wang (2020a); Wang et al. (2017), which are essential for applications like legacy software migration and security instrumentation tasks Wong et al. (2023); Dinesh et al. (2020).
@@
Figure 1 illustrates the transformation from the source C code to a binary file, assembly code (ASM), and pseudo-code decompiled from Ghidra. In this pseudo-code, the original nested for structure is replaced with a less intuitive combination of a do-while loop inside another while loop. Furthermore, array indexing like num[i] is decompiled into complicated pointer arithmetic such as *(float *)(param_2 + (long)local_24 * 4). The decompiled output also exhibits syntactical errors, with the function return type being converted to undefined4. Overall, traditional decompilation tools often strip away the syntactic clarity provided by high-level languages and do not ensure the correctness of syntax, posing significant challenges even for skilled developers to reconstruct the algorithmic logic Wong et al. (2023); Hu et al. (2024).
@@
To address the limitations of previous studies, we propose LLM4Decompile, the first and largest open-source LLM series with sizes ranging from 1.3B to 33B parameters specifically trained to decompile binary code. To the best of our knowledge, there’s no previous study attempts to improve the capability of LLM-based decompilation in such depth or incorporate such large-scale LLMs. Based on the End2end-Decompile approach, we introduce three critical steps: data augmentation, data cleaning, and two-stage training, to optimize the LLM training process and introduce the LLM4Decompile-End models to decompile binary directly. Specifically, our LLM4Decompile-End-6.7B model demonstrates a successful decompilation rate of 45.4% on HumanEval Chen et al. (2021) and 18.0% on ExeBench Armengol-Estapé et al. (2022), far exceeding Ghidra Ghidra (2024a) or GPT-4o OpenAI (2023) by over 100%. Additionally, we improve the Refined-Decompile strategy by examining the efficiency of Ghidra’s decompilation process, augmenting and filtering data to fine-tune the LLM4Decompile-Ref models, which excel at refining Ghidra’s output. Experiments suggest a higher performance ceiling for the enhanced Refined-Decompile approach, with 16.2% improvement over LLM4Decompile-End. Additionally, we assess the risks associated with the potential misuse of our model under obfuscation conditions commonly used in software protection. Our findings indicate that neither our approach nor Ghidra can effectively decompile obfuscated code, mitigating concerns about unauthorized use for infringement of intellectual property.
@@
  • We optimize the LLM training process and introduce LLM4Decompile-End models, which set a new performance standard of direct binary decompilation, significantly surpassing GPT-4o and Ghidra by over 100% in terms of re-executability on the HumanEval and ExeBench benchmarks.
  • We improve the Refined-Decompile approach to fine-tune the LLM4Decompile-Ref models, enabling them to effectively refine the decompiled results from Ghidra and achieve further 16.2% re-executability enhancements over LLM4Decompile-End.
@@ 2 Related Work
The practice of reversing executable binaries to their source code form, known as decompilation, has been researched for decades Miecznikowski and Hendren (2002); Nolan (2012); Katz et al. (2019). Traditional decompilation relies on analyzing the control and data flows of program Brumley et al. (2013), and employing pattern matching, as seen in tools like Hex-Rays Ida pro Hex-Rays (2024) and Ghidra Ghidra (2024a). These systems attempt to identify patterns within a program’s control-flow graph (CFG) that corresponding to standard programming constructs such as conditional statements or loops. However, the output from such decompilation processes tends to be a source-code-like representation of assembly code, including direct translations of variables to registers, use of gotos, and other low-level operations instead of the original high-level language constructs. This output, while often functionally similar to the original code, is difficult to understand and may not be re-executable Liu and Wang (2020b); Wong et al. (2023). Drawing inspiration from neural machine translation, researchers have reformulated decompilation as a translation exercise, converting machine-level instructions into readable source code Katz et al. (2019). Initial attempts in this area utilized recurrent neural networks (RNNs) Katz et al. (2018) for decompilation, complemented by error-correction techniques to enhance the outcomes.
@@
Motivated by the success of Large Language Models Li et al. (2023); Rozière et al. (2023); Guo et al. (2024), researchers have employed LLMs for decompilation, primarily through two approaches—Refined-Decompile and End2end-Decompile. In particular, Refined-Decompile prompts the LLMs to refine results from traditional decompilation tools like Ghidra or IDA Pro. For instance, DeGPT Hu et al. (2024) enhances Ghidra’s readability by reducing cognitive load by 24.4%, while DecGPT Wong et al. (2023) increases IDA Pro’s re-executability rate to over 75% by integrating error messages into its refinement process. These approaches, however, largely ignore the fact that LLMs are designed primarily for high-level programming languages Li et al. (2023); Rozière et al. (2023); Guo et al. (2024), and their effectiveness with binary files is not well-established. End2end-Decompile, on the other hand, fine-tunes LLMs to decompile binaries directly. Early open-source models like BTC Hosseini and Dolan-Gavitt (2022) and recent development Slade Armengol-Estapé et al. (2023) adopt the language model with around 200 million parameters Lewis et al. (2020a) to fine-tune for decompilation. While Nova Jiang et al. (2023), which is not open-sourced, develops a binary LLM with 1 billion parameters and fine-tunes it for decompilation. Consequently, the largest open-source model in this domain is limited to 200M. Whereas utilizing larger models trained on broader datasets has proven to substantially improve the performance Hoffmann et al. (2024); Kaplan et al. (2020); Rozière et al. (2023).
@@ 3.1.2 Optimize LLM4Decompile-End
Our final step for training optimization aims to educate the model with binary knowledge, and includes two-stage training. In the first stage, we train the model with a large corpus of compilable but not linkable (executable) data. Note that it’s significantly easier to extract C code that is compilable but not linkable da Silva et al. (2021); Armengol-Estapé et al. (2022). Such not-executable binary object code will closely resemble its executable version except it lacks linked addresses for external symbols. Therefore, in the first stage, we use the extensive compilable codes to ground our model in binary knowledge. In the second stage, we refine the model using executable code to ensure its practical applicability. We also conduct an ablation study for the two-stage training in Section 4.1.2. Comparison between compilable and executable data is detailed in Appendix B.
@@ 3.2.2 Refine Ghidra by LLM4Decompile-Ref
Decompiling the executable code with Ghidra (Figure 3) is time-consuming due to the complex nature of the executables in ExeBench, which include numerous external functions and IO wrappers. Ghidra Headless Ghidra (2024b) requires 2 seconds per sample using 128-core multiprocessing. Given such a high computational load, and the high similarities between non-executable and executable binaries, we choose to decompile the non-executable files using Ghidra. This choice significantly reduces the time to 0.2 seconds per sample, enabling us to efficiently gather large amounts of training data.
@@ 4.1.1 Experimental Setups
To evaluate the models, we introduce HumanEval Chen et al. (2021) and ExeBench Armengol-Estapé et al. (2022) benchmarks. HumanEval is the leading benchmark for code generation assessment and includes 164 programming challenges with accompanying Python solutions and assertions. We converted these Python solutions and assertions into C, making sure that they can be compiled with the GCC compiler using standard C libraries Free Software Foundation (2024) and pass all the assertions, and name it HumanEval-Decompile. ExeBench consists of 5000 real-world C functions taken from GitHub with IO examples. Note that the HumanEval-Decompile consists of individual functions that depend only on the standard C library. In contrast, ExeBench includes functions extracted from real-world projects with user-defined structures and functions [^2] Exebench provides comparison of the test set against the GitHub population using nine distinct code complexity metrics, confirming that the characteristics of the testing functions are aligned with functions in larger real-world projects..
@@
Model/Benchmark HumanEval-Decompile ExeBench O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG DeepSeek-Coder-6.7B 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 GPT-4o 30.49 11.59 10.37 11.59 16.01 4.43 3.28 3.97 3.43 3.78 LLM4Decompile-End-1.3B 47.20 20.61 21.22 20.24 27.32 17.86 13.62 13.20 13.28 14.49 LLM4Decompile-End-6.7B 68.05 39.51 36.71 37.20 45.37 22.89 16.60 16.18 16.25 17.98 LLM4Decompile-End-33B 51.68 25.56 24.15 24.75 31.54 18.86 14.65 13.96 14.11 15.40
@@
Model/Benchmark HumanEval-Decompile ExeBench O0 O1 O2 O3 AVG O0 O1 O2 O3 AVG Compilable-1.3B 42.68 16.46 16.46 17.07 23.17 5.68 4.46 4.16 4.43 4.68 Compilable-6.7B 51.83 33.54 32.32 32.32 37.50 7.52 6.49 6.71 6.60 6.83 Executable-1.3B 19.51 12.80 12.80 11.59 14.18 21.94 19.46 19.31 19.50 20.05 Executable-6.7B 37.20 18.29 22.56 17.07 23.78 29.38 25.98 25.91 25.49 26.69
@@
We selected two key baselines for comparison. First, GPT-4o OpenAI (2023) represents the most capable LLMs, providing an upper bound on LLM performance. Second, DeepSeek-Coder Guo et al. (2024) is selected as the current SOTA open-source Code LLM. It represents the forefront of publicly available models specifically tailored for coding tasks. While recent work Slade Armengol-Estapé et al. (2023) fine-tunes language model for decompilation, it relies on intermediate compiler outputs, specifically, the *.s files. In practice, however, such intermediate files are rarely released by developers. Therefore, we focus on a more realistic approach, and consider decompilation only from the binaries, for further discussions please refer to Appendix B.
@@ 4.1.2 Experimental Results
Table 3: Main comparison of Refined-Decompile approaches for re-executability rate and Edit Similarity on HumanEval-Decompile benchmark. “+GPT-4o” refers to enhance the Ghidra results with GPT-4o, “+LLM4Decompile-Ref” means refining Ghidra results with the fine-tuned LLM4Decompile-Ref models. Note that the 33B model was trained using only 200M tokens, which is just 10% of the tokens used for the 1.3B/6.7B/22B model. For the 22B model, Please refer to Appendix D.
@@
As discussed in Section 4.1.1, our training data comprises two distinct sets: 7.2 million compilable functions (non-executable) and 1.6M executable functions. We conducted an ablation study using these datasets, and the results are displayed in Table 2. Here, “Compilable” denotes the model trained solely on compilable data, while “Executable” indicates models trained exclusively on executable data. Notably, the binary object from compilable functions lacks links to function calls, which is similar in text distribution to the HumanEval-Decompile data, consisting of single functions dependent only on standard C libraries. Consequently, the 6.7B model trained only on compilable data successfully decompiled 37.5% of HumanEval-Decompile functions, but only 6.8% on ExeBench, which features real functions with extensive user-defined functions. On the other hand, the 6.7B model trained solely on executable data achieved a 26.7% re-executability rate on the ExeBench test set but faced challenges with single functions, with only a 23.8% success rate on HumanEval-Decompile due to the smaller size of the training corpus. Limited by the space, we present further analysis in Appendix C.